13. Two point charges are brought closer together, increasing the force between them by a factor of 25. By what factor was their separation decreased?
Answer #1

force is inversely proportional to the separation distance squared.

F1 = kq1q2 / r1^2

F2 = kq1q2 / r2^2

ratio of the forces we can determine the ratio of the separation distances

F1/F2 = (r1/r2)^2

r2/r1 = sqrt(F1/F2)

r2/r1 = sqrt(1/25)

r2/r1 = 1/5

The separation decreased by a factor of 5